please help me with this question:
The area of a rectange is 338 mm squared. the height is thirteen less than three times teh base. What is the perimeter?

Answer by checkley75(3666)

338=b*h where h=3b-13
338=b(3b-13)
338=3b^2-13b
3b^2-13b-338=0
(3b+26)(b-13)=0
b-13=0
b=13 answer for the base.
h=3*13-13
h=39-13
h=26 answer for the height.
perimeter=2*b+2*h
p=2*13+2*26
p=26+52
p=78 answer for the perimeter.